# Hybrid Many-Objective Optimization in Probabilistic Mission Design for Compliant and Effective UAV Routing

### v1

@@ Abstract.
Advanced Aerial Mobility encompasses many outstanding applications that promise to revolutionize modern logistics and pave the way for various public services and industry uses. However, throughout its history, the development of such systems has been impeded by the complexity of legal restrictions and physical constraints. While airspaces are often tightly shaped by various legal requirements, Unmanned Aerial Vehicles (UAV) must simultaneously consider, among others, energy demands, signal quality, and noise pollution. In this work, we address this challenge by presenting a novel architecture that integrates methods of Probabilistic Mission Design (ProMis) (Kohaut et al., 2023a, 2024) and Many-Objective Optimization (Hohmann et al., 2023) for UAV routing. Hereby, our framework is able to comply with legal requirements under uncertainty while producing effective paths that minimize various physical costs a UAV needs to consider when traversing human-inhabited spaces. To this end, we combine hybrid probabilistic first-order logic for spatial reasoning with mixed deterministic-stochastic route optimization, incorporating physical objectives such as energy consumption and radio interference with a logical, probabilistic model of legal requirements. We demonstrate the versatility and advantages of our system in a large-scale empirical evaluation over real-world, crowd-sourced data from a map extract from the city of Paris, France, showing how a network of effective and compliant paths can be formed.
@@ 1. Introduction
Unmanned Aerial Vehicles (UAV) have emerged as a novel form of autonomous transportation systems, with numerous applications in logistics, public, and industry promising to disrupt the market and revolutionize the respective fields. However, their employment depends on two factors: First, the availability of robust routing schemes that can satisfy many physical objectives at once, such as energy and time requirements, signal strengths, or noise pollution in human-inhabited spaces. Second, the ability to delicately navigate airspace regulations, such as those created for Unmanned Traffic Management (UTM) or adopted by local authorities. In combination, providing a UAV routing scheme that allows both of these criteria to be factored in is an outstanding challenge in enabling Advanced Air Mobility (AAM) applications.
@@
With regard to physical objectives, prior work has demonstrated the application of hybrid many-objective optimization techniques (Hohmann et al., 2023). By combining exact graph-based search with a metaheuristic evolutionary optimization, a set of smooth paths that individually trade-off the pre-defined objectives is obtained. Although this enables navigating within physical optimization goals with fine-granular control over their relative importance, modeling and adhering to complex sets of traffic rules remained a challenge. To this end, Probabilistic Mission Design (ProMis) has been presented to encode operational constraints such as UTM rules, manufacturer restrictions, and operator preferences as hybrid probabilistic first-order logic programs (Kohaut et al., 2023b). Hence, this framework allows for probabilistic inference across the agent’s navigation space, resulting in scalar fields of independent, identically distributed (i.i.d.) probabilities of all modeled laws being satisfied under uncertainty in the environment. Hence, mission design, as they are crucial for enabling AAM applications such as deciding clearance, explaining the decision, and proposing better mission settings, can be implemented over ProMis (Kohaut et al., 2024).
@@
In this work, we combine these approaches into a single system for UAV routing under uncertainty, allowing AAM scenarios to take place in an effective manner while complying with the local legal restrictions established in human-inhabited spaces through probabilistic and symbolic reasoning. Given start- and end-points, our framework produces a Pareto-set of physically effective and legally compliant routes from which the agent can select according to its individual weighting of their respective importance. Furthermore, clearance for the resulting paths can be decided independently, allowing one to discard any potential candidates that do not sufficiently respect regulations. To this end, we demonstrate how the impact of high-level parameters controlling the mission’s overall setting, such as the employed license of the pilot or time-of-day, can be explained to the UAV operator and automatically selected for optimal rule adherence.
@@
We present Probabilistic Mission Design (ProMis) with Hybrid Many-Objective Optimization for UAV routing to consider physical objectives and legal constraints in urban AAM jointly.
We introduce the Constitutional Language (CoLa), a novel description language for encoding an agent’s objectives across its state space and paths to combine legal constraints and physical objectives in a unified fashion. To this end, CoLa makes use of spatial reasoning under uncertainty facilitated via Statistical Relational Maps (StaR Maps) (Flade et al., 2024).
We demonstrate how our system can decide Clearance for a given mission, Explain the impact of high-level, semantic parameters, and Optimize for ideal compliance.
We publish our implementations of (i) our Versatile Intelligent Aerial Streets (VIAS111https://github.com/NikHoh/VIAS) algorithm for 3D path-planning under many objectives and (ii) our overarching system for Probabilistic Mission Design and Routing222https://www.github.com/HRI-EU/ProMis as presented in this work.
We proceed as follows.
As our work is synthesized from a background of probabilistic environment representations, many-objective routing, and neuro-symbolic reasoning systems, we discuss each of these areas and touch upon recent developments of regulatory frameworks for UAVs in Section 2.
With this overview in mind, we present our methods in Section 3. Starting with the architecture of our approach, made up of StaR Maps, ProMis, and the VIAS many-objective optimizer, we present the individual building blocks in detail, including the objectives we consider in this work and the syntax and semantics of CoLa.
To underpin our findings, we present an exhaustive experimental study over real-world data from the city of Paris in Section 4, showing how our system is able to facilitate networks of physically effective and legally compliant paths for UAVs even in complex urban environments.
Finally, we conclude with a summary of this work, discuss its limitations, and point toward future work in Section 5.
@@ 2. Related Work
In Europe, the Single European Sky Air Traffic Management Research (SESAR) initiative has significantly influenced the advancement of Unmanned Aerial Vehicle (UAV) applications and Urban Air Mobility (UAM). However, with subsequent editions of the SESAR Master Plan, particularly the 2015 release (SESAR Joint Undertaking, 2015) and the 2017 Drone Outlook Study (SESAR Joint Undertaking, 2017), the integration of UAV into European airspace gained momentum.
Regulations and operational restrictions have been introduced concurrently to ensure the safety of drone operations. Notably, the EU regulation 2019/947 (The European Commission, 2019) and Specified Operations Risk Assessment concepts (European Union Aviation Safety Agency, 2019) underscores the coexistence of unmanned and manned aircraft within shared airspace under rigorous safety checks.
@@ 2.1. Representation
Numerous sensors can be employed by the UAV itself, specialized mapping platforms, or personal computing devices for crowd-sourcing data to gather the necessary information on the environment. They range from proprioceptive sensors, e.g., using the Global Navigation Satellite System (Joubert et al., 2020) and Inertial Measurement Units (Zhang et al., 2012; Gao et al., 2022) which inform about the agent’s own location, to exteroceptive sensors such as cameras (Alkendi et al., 2021; Flade et al., 2018), Lidar (Demir and Fujimura, 2019; Petrlik et al., 2021), and Radar (Ward and Folkesson, 2016; Michalczyk et al., 2022) which inform about the environment.
Frequently, maps created from optical data (Levinson and Thrun, 2010; Huang et al., 2019; Lategahn and Stiller, 2012) may lack deeper semantics and are unsuitable for navigation beyond simple obstacle avoidance, e.g., unstructured point clouds or unlabeled structures from image data. To remedy this issue, vision-based classification approaches such as segment anything (Kirillov et al., 2023), and object-centric vision (Locatello et al., 2020; Kipf et al., 2022) can be employed to enrich the map with semantic labels for more principled navigation approaches. Of course, human insights might also be employed to label the map data.
@@
Recently, we have presented the Statistical Relational Map (StaR Map) as a method for encoding uncertain environments in a relational fashion (Flade et al., 2024). Rather than storing geometric information, StaR Maps represent uncertain environments through a collection of categorical and continuous parametric distributions, each associated with a spatial relation and a category of geographic features. Hereby, integration with probabilistic and logical reasoning frameworks is enabled and discussed further as part of our architecture in Section 3.
@@ 2.2. Routing
Given sufficient knowledge about the agent’s environment, decisions ought to be made on how to travel from the current position to a desired goal. Although achieving the shortest time to fulfill a task is often the obvious target, many objectives may be imposed on the path planning.
@@
In the first instance, an agent may consider a single objective to decide a suitable path. For example, risk minimization of collisions may be employed to decide velocity profiles that avoid getting too close to other traffic participants even under uncertainty (Damerow and Eggert, 2014). Similarly, given deeper knowledge of the environment, such approaches can be enhanced by considering where other agents’ paths may lead, allowing for more decisive and safe planning (Puphal and Eggert, 2023).
@@
Since no singular optimal solution exists when considering multiple objectives simultaneously, a Pareto set of paths may be produced representing a range of trade-offs between the individual cost functions. One early approach for many-objective path planning was proposed by Nikolos et al. (Nikolos et al., 2003), who optimize UAV paths for flights over rough terrain. In a later approach, Rubio-Hervas et al. (Rubio-Hervas et al., 2018) adopted an urban setting in Singapore to plan 3D paths for UAVs. They propose a special path representation composed of the distance and the angle between waypoints and a straight line connecting the path’s start and goal. Evolutionary algorithms have been shown to successfully reduce the computational load of solving many-objective path planning, e.g., as in work by Ghambari et al. (Ghambari et al., 2020). To optimize a path regarding its length and risk, they employ the NSGA-II algorithm (Deb et al., 2002), a state-of-the-art evolutionary algorithm for multi-objective optimization capable of calculating a Pareto set. Because using a grid-based path representation can result in sharp turns and inefficient jagged paths, so a smooth spline curve representation is often preferred. For a deeper dive into the topic, we refer the interested reader to an introduction on multi-objective optimization (Emmerich and Deutz, 2018).
@@
Our work is based on two core ingredients for UAV routing, namely hybrid many-objective optimization for 3D path planning (Hohmann et al., 2023) and the Probabilistic Mission Design framework (Kohaut et al., 2023a). While the former is based on a graph-based search and a subsequent evolutionary search for fine-granular routing, the latter shows how reasoning on legal frameworks and operational preferences can be enabled via hybrid probabilistic logic programs, producing independent and identically distributed probabilities of the requested model being satisfied at each continuous point in state-space.
@@
By combining these two frameworks, a powerful UAV routing pipeline emerges that consumes a hybrid probabilistic, logical representation of legal constraints alongside physical objectives to produce a Pareto-set of proposal paths. Furthermore, we demonstrate how mission design tasks such as Clearance, Explanation, and Optimization of proposal path and mission parameters are supported (Kohaut et al., 2024), forming an interpretable, adaptable, and safe navigation system.
@@ 2.3. Reasoning
When creating autonomous mobile agents, it is necessary to consider not only their ability to reach their goal without collisions but also to reason on the circumstances to conclude safe and compliant decisions. Endowing an agent with such capabilities, e.g., ensuring that local legal regulations are satisfied, is subject to ongoing research.
@@
Here, we build on the Probabilistic Mission Design (ProMis) framework (Kohaut et al., 2023b) and its capabilities for facilitating crucial tasks such as mission Clearance, Explanation, and Optimization (Kohaut et al., 2024). It aligns with preceding research efforts, offering a formal symbolic approach to probabilistic classification and mission design through spatial reasoning based on systems such as StaR Maps (Flade et al., 2024). Notably, ProMis provides an adaptable and interpretable interface to mission design while allowing arbitrary queries such as on the satisfaction of local airspace regulations, thus facilitating comprehensive compliance assessment and decision-making. Moreover, ProMis accommodates evolving regulatory frameworks and operational priorities within its probabilistic framework, such as battery lifetime (An et al., 2023) or weather-effects (Schuchardt et al., 2022).
@@ 3.1. Architecture
We begin by outlining our architecture for compliant and effective Unmanned Aerial Vehicles (UAV) routing. It can be divided into the three pillars of environment representation, mission design, and UAV routing through many-objective optimization. We base each of these pillars on our prior publications on Statistical Relational Maps (StaR Maps) (Flade et al., 2024), Probabilistic Mission Design (ProMis) with Clearance, Explanation and Optimization techniques (Kohaut et al., 2023a, 2024) and Hybrid Mani-Objective Optimization for 3D UAV routing (Hohmann et al., 2023) respectively. The resulting architecture for such effective and compliant UAV routing is illustrated in Figure 1.
We assume a UAV with access to geographic data such as the local road network, buildings, and infrastructure, as well as (neural) sensors for online perception. Hence, the agent can compute and maintain a StaR Map (see Section 3.2) to represent its environment in a hybrid probabilistic set of spatial relations.
Said StaR Map is employed to parameterize the agent’s Constitution (see Section 3.4), i.e., a unified description of the applicable local airspace regulations and physical objectives to be considered during travel. While physical objectives are provided in analytical models (see Section 3.3), inference in probabilistic first-order logic estimates how likely states across the navigation space will satisfy regulations.
Based on these objectives, we employ the Versatile Intelligent Aerial Streets (VIAS) routing algorithm, resulting in a Pareto-front of paths (see Section 3.5). Given a weighting of the objectives, a proposed path is decided and checked for Clearance based on the traffic restrictions encoded in the Constitution (see Section 3.6). If Clearance is granted, the agent accepts and executes the path. Otherwise, an Explanation can be given to understand which mission parameters impacted the decision on Clearance the most (see Section 3.7), and the Constitution’s parameters may be searched for a more suitable setting, e.g., changing at what time of day the mission shall be carried out (see Section 3.8).

### v2

@@ Abstract.
Advanced Aerial Mobility encompasses many outstanding applications that promise to revolutionize modern logistics and pave the way for various public services and industry uses. However, throughout its history, the development of such systems has been impeded by the complexity of legal restrictions and physical constraints. While airspaces are often tightly shaped by various legal requirements, Unmanned Aerial Vehicles (UAV) must simultaneously consider, among others, energy demands, signal quality, and noise pollution. In this work, we address this challenge by presenting a novel architecture that integrates methods of Probabilistic Mission Design (ProMis) (Kohaut et al., 2023a, 2024) and Many-Objective Optimization (Hohmann et al., 2023) for UAV routing. Hereby, our framework facilitates compliance with legal requirements under uncertainty while producing effective paths that minimize various physical costs a UAV needs to consider when traversing human-inhabited spaces. To this end, we combine hybrid probabilistic first-order logic for spatial reasoning with mixed deterministic-stochastic route optimization, incorporating physical objectives such as energy consumption and radio interference with a logical, probabilistic model of legal requirements. We demonstrate the versatility and advantages of our system in a large-scale empirical evaluation over real-world, crowd-sourced data from a map extract from the city of Paris, France, showing how a network of effective and compliant paths can be formed.
@@ 1. Introduction
Unmanned Aerial Vehicles (UAV) have emerged as a novel form of autonomous transportation systems, with numerous applications in logistics, public, and industry, promising to disrupt the market and revolutionize the respective fields. However, their employment depends on two factors: First, the availability of robust routing schemes that can satisfy many physical objectives simultaneously, such as energy and time requirements, signal strengths, or noise pollution in human-inhabited spaces. Second, the ability to delicately navigate airspace regulations, such as those created for Unmanned Traffic Management (UTM) or adopted by local authorities. Providing a UAV routing scheme that allows both criteria to be factored in is an outstanding challenge in enabling Advanced Air Mobility (AAM) applications.
@@
With regard to physical objectives, prior work has demonstrated the application of hybrid many-objective optimization techniques (Hohmann et al., 2023). That is, by combining exact graph-based search with a metaheuristic evolutionary optimization, a set of smooth paths that individually trade off the pre-defined objectives can be obtained. Although this enables navigating within physical optimization goals with fine-granular control over their relative importance, modeling and adhering to complex traffic rules remains challenging. To this end, Probabilistic Mission Design (ProMis) has been presented to encode operational constraints such as UTM rules, manufacturer restrictions, and operator preferences as hybrid probabilistic first-order logic programs (Kohaut et al., 2023b). Hence, ProMis allows for probabilistic inference across the agent’s navigation space, resulting in scalar fields of independent, identically distributed (i.i.d.) probabilities of all modeled laws being satisfied under uncertainty in the environment. Mission design, as it is crucial for enabling AAM applications such as deciding clearance, explaining the decision, and proposing better mission settings, can be implemented over ProMis (Kohaut et al., 2024).
@@
In this work, we combine these approaches into a single system for UAV routing under uncertainty, allowing AAM scenarios to take place in an effective manner while complying with the local legal restrictions established in human-inhabited spaces through probabilistic and symbolic reasoning. Given start- and end-points, our framework produces a Pareto set of physically effective and legally compliant routes from which the agent can select according to its individual weighting of their respective importance. Furthermore, clearance for the resulting paths can be decided independently, allowing one to discard any potential candidates that do not sufficiently respect regulations. To this end, we demonstrate how the impact of high-level parameters controlling the mission’s overall setting, such as the employed license of the pilot or time-of-day, can be explained to the UAV operator and automatically selected for optimal rule adherence.
@@
We present Probabilistic Mission Design (ProMis) with Hybrid Many-Objective Optimization for UAV routing to jointly consider physical objectives and legal constraints in urban AAM.
We introduce the Constitutional Language (CoLa), a novel description language for encoding an agent’s objectives across its state space and paths to combine legal constraints and physical objectives in a unified fashion. To this end, CoLa uses spatial reasoning under uncertainty facilitated via Statistical Relational Maps (StaR Maps) (Flade et al., 2024).
We demonstrate how our system can decide Clearance for a given mission, Explain the impact of high-level semantic parameters, and Optimize for ideal compliance.
We publish our implementations of (i) our Versatile Intelligent Aerial Streets (VIAS111github.com/NikHoh/VIAS) algorithm for 3D path-planning under many objectives and (ii) our overarching system for Probabilistic Mission Design and Routing222github.com/HRI-EU/ProMis as presented in this work.
The UAV routing framework presented in this paper builds on and synthesizes several prior works. Specifically, we integrate StaR Maps (Flade et al., 2024) as a hybrid probabilistic environment model, and parameterize a probabilistic logic representation of UAV behavior, similar to ProMis (Kohaut et al., 2023b), to enable high-level decision-making during mission execution (Kohaut et al., 2024). We expand on these ideas with our CoLa language, which introduces a description layer that unifies social and physical objectives for effective and compliant UAV routing. Ultimately, this setup guides the path planning process within our hybrid many-objective optimization framework, VIAS (Hohmann et al., 2023), allowing us to apply the aforementioned models in a large-scale, 3D routing evaluation employing real-world data.
In the following, we discuss the backgrounds of probabilistic environment representations, many-objective routing, and neuro-symbolic reasoning, and touch upon recent developments of regulatory frameworks for UAVs in Section 2.
With this overview in mind, we present our methods in Section 3. Starting with the architecture of our approach, made up of StaR Maps, ProMis, and the VIAS many-objective optimizer, we lay out our framework’s building blocks in detail, including the objectives we consider in this work and the syntax and semantics of CoLa.
Our findings are underpinned by an exhaustive experimental study over crowd-sourced, real-world data from Paris, France, in Section 4. Hence, we show how our system yields networks of physically effective and legally compliant paths for UAVs, even in complex urban environments.
Finally, we summarize our work, discuss its limitations, and point toward future work in Section 5.
@@ 2. Related Work
When discussing novel approaches to compliance in intelligent transportation systems, it is important to consider the ongoing development of public regulations. In Europe, the Single European Sky Air Traffic Management Research (SESAR) initiative has significantly influenced the advancement of Unmanned Aerial Vehicle (UAV) applications and Urban Air Mobility (UAM). With subsequent editions of the SESAR Master Plan, particularly the 2015 release (SESAR Joint Undertaking, 2015) and the 2017 Drone Outlook Study (SESAR Joint Undertaking, 2017), the integration of UAVs into the European airspace gained momentum.
Regulations and operational restrictions have been introduced concurrently to ensure the safety of drone operations. Notably, the EU regulation 2019/947 (The European Commission, 2019) and Specified Operations Risk Assessment concepts (European Union Aviation Safety Agency, 2019) underscore the coexistence of unmanned and manned aircraft within shared airspace under rigorous safety checks.
@@ 2.1. Representation
Numerous sensors can be employed by the UAV, specialized mapping platforms, or personal computing devices for crowd-sourcing data to gather the necessary environmental information. They range from proprioceptive sensors, e.g., using the Global Navigation Satellite System (Joubert et al., 2020) and Inertial Measurement Units (Zhang et al., 2012; Gao et al., 2022) which inform about the agent’s own location, to exteroceptive sensors such as cameras (Alkendi et al., 2021; Flade et al., 2018), Lidar (Demir and Fujimura, 2019; Petrlik et al., 2021), and Radar (Ward and Folkesson, 2016; Michalczyk et al., 2022) which inform about the environment.
Frequently, maps created, e.g., from optical sensors (Levinson and Thrun, 2010; Huang et al., 2019; Lategahn and Stiller, 2012) may lack deeper semantics and are unsuitable for navigation beyond simple obstacle avoidance, e.g., unstructured point clouds or unlabeled structures from image data. To remedy this issue, vision-based classification approaches such as segment anything (Kirillov et al., 2023) and object-centric vision (Locatello et al., 2020; Kipf et al., 2022) can be employed to enrich the map with semantic labels for more principled navigation approaches. Of course, human insights might also be employed to label the mapped environment features.
@@
Recently, we have presented the Statistical Relational Map (StaR Map) as a method for encoding uncertain environments in a relational fashion (Flade et al., 2024). Rather than storing geometric information, StaR Maps consume data sources as discussed above to represent uncertain environments through sets of categorical and continuous parametric distributions, each associated with a spatial relation and a set of geographic features. Hence, StaR Maps allow for explicit consideration of the inaccuracies of the mapping process and employed sensors with the annotated semantics of the respective environment features. Furthermore, they can be compiled into the parameters of probabilistic and logical reasoning frameworks such as our Constitutional Language (CoLa) as discussed in Section 3.
@@ 2.2. Routing
Given sufficient knowledge about the agent’s environment, decisions ought to be made on how to travel from the current position to a desired goal. Although achieving the shortest time to fulfill a task is often the obvious target, many objectives may be imposed on the agent’s path planning.
@@
First, an agent may consider a single objective to decide a suitable path. For example, risk minimization of collisions may be employed to decide velocity profiles that avoid getting too close to other traffic participants even under uncertainty (Damerow and Eggert, 2014). Similarly, given more profound knowledge of the environment, such approaches can be enhanced by considering other agents’ paths, allowing for more decisive and safe planning (Puphal and Eggert, 2023).
@@
Since no singular optimal solution exists when considering multiple objectives simultaneously, a Pareto set of paths may be produced representing a range of trade-offs between the individual cost functions. One early approach for many-objective path planning optimizes UAV paths for flights over rough terrain (Nikolos et al., 2003). In a later approach, an urban setting in Singapore was adopted to plan 3D paths for UAVs (Rubio-Hervas et al., 2018). A special path representation was proposed, which was composed of the distance and the angle between waypoints and a straight line connecting the path’s start and goal.
Evolutionary algorithms have been shown to successfully reduce the computational load of solving many-objective path planning (Ghambari et al., 2020). To optimize a path regarding its length and risk, the NSGA-II algorithm (Deb et al., 2002) was employed, a state-of-the-art evolutionary algorithm for multi-objective optimization capable of calculating a Pareto set. Because using a grid-based path representation can result in sharp turns and inefficient jagged paths, a smooth spline curve representation is often preferred. For a deeper dive into the topic, we refer the interested reader to an introduction on multi-objective optimization (Emmerich and Deutz, 2018).
@@
Our work is based on two core UAV routing ingredients: hybrid many-objective optimization for 3D path planning (Hohmann et al., 2023) and the Probabilistic Mission Design (ProMis) framework (Kohaut et al., 2023a). While the former builds on a graph-based search and a subsequent evolutionary search for fine-granular routing, the latter shows how reasoning on legal frameworks and operational preferences can be enabled via hybrid probabilistic logic programs, producing independent and identically distributed probabilities of the requested model being satisfied at each continuous point in state space.
@@
By combining these two frameworks, a powerful UAV routing pipeline emerges that consumes a hybrid probabilistic, logical representation of legal constraints alongside physical objectives to produce a Pareto set of proposal paths. This not only yields compliant and effective paths but also allows for individual trade-offs between objectives without the need to enforce each with the same strictness. Furthermore, we demonstrate how mission design tasks such as Clearance, Explanation, and Optimization of proposal path and mission parameters are supported (Kohaut et al., 2024), forming an interpretable, adaptable, and safe navigation system.
@@ 2.3. Reasoning
When creating autonomous mobile agents, it is necessary to consider their ability to reach their goal without collisions and simultaneously reason on the circumstances to conclude safe and compliant decisions. Endowing an agent with such capabilities, e.g., ensuring that local legal regulations are satisfied, is subject to ongoing research.
@@
In this work, we build on ProMis and its capabilities for facilitating crucial tasks such as mission Clearance, Explanation, and Optimization (Kohaut et al., 2024). It aligns with preceding research efforts, offering a formal symbolic approach to probabilistic classification and mission design through spatial reasoning based on systems such as StaR Maps (Flade et al., 2024). Notably, ProMis provides an adaptable and interpretable interface to mission design while allowing arbitrary queries such as those about the satisfaction of local airspace regulations, thus facilitating comprehensive compliance assessment and decision-making. Moreover, ProMis accommodates evolving regulatory demands and operational priorities within its probabilistic framework, such as battery lifetime (An et al., 2023) or weather effects (Schuchardt et al., 2022). By extending ProMis with CoLa as a high-level modeling language, we move beyond its original focus on compliance and introduce physical objectives into the model, allowing ProMis’ application for real-world UAV routing with path effectiveness in mind.
@@ 3.1. Architecture
We begin by outlining our architecture for compliant and effective Unmanned Aerial Vehicles (UAV) routing. It can be divided into the three pillars of environment representation, mission design, and UAV routing through many-objective optimization. We base each of these pillars on our prior publications on Statistical Relational Maps (StaR Maps) (Flade et al., 2024), Probabilistic Mission Design (ProMis) with Clearance, Explanation and Optimization techniques (Kohaut et al., 2023a, 2024), and Hybrid Many-Objective Optimization for 3D UAV routing (Hohmann et al., 2023) respectively. The resulting architecture for such effective and compliant UAV routing is illustrated in Figure 1.
We assume a UAV with access to data from a Geographic Information System, such as the local road network, buildings, and infrastructure, as well as (neural) sensors for online perception of more dynamic environment features such as people or other vehicles. With this information at hand, the agent can compute and maintain a StaR Map (see Section 3.2) to represent its environment in a hybrid probabilistic set of spatial relations between the agent’s state space and the environment, e.g., probabilities wether the agent would be over a certain feature or at a required distance in a specific state.
The StaR Map is employed to parameterize the agent’s Constitution (see Section 3.4), i.e., a unified description of the applicable local airspace regulations and physical objectives to be considered during travel. While physical objectives are provided in analytical models (see Section 3.3), inference in probabilistic first-order logic allows the agent to estimate how likely states across the navigation space will satisfy regulations.
Based on these objectives, we employ the Versatile Intelligent Aerial Streets (VIAS) routing algorithm, resulting in a Pareto front of paths (see Section 3.5). Given a weighting of the objectives, a proposed path is decided and checked for Clearance based on the traffic restrictions encoded in the Constitution (see Section 3.6). If Clearance is granted, the agent accepts and executes the path. Otherwise, an Explanation can be given to understand which mission parameters impacted the decision on Clearance the most (see Section 3.7), and the Constitution’s parameters may be searched for a more suitable setting, e.g., changing at what time of day the mission shall be carried out (see Section 3.8).